OpenStudy (aurorab):

Jake is comparing the prices of two mattress cleaning companies. Company A charges $30 per mattress and an additional $13 as service charges. Company B charges $28 per mattress and an additional $15 as service charges. Part A: Write equations to represent Company A and Company B's total mattress cleaning charges for a certain number of mattresses. Define the variable used in the equations. Part B: Which company would charge less for cleaning 4 mattresses? Justify your answer. Part C: How much money is saved by using the services of Company B instead of Company A to clean 7 mattresses?

OpenStudy (aurorab):

HELPP!!!!!!! @skullpatrol @sisco0123 @dan815

OpenStudy (shivaniits):

let the no of mattresses are x then for company A : \[t_a=30x+13\] here \[t_a\] for total cleaning charges for company A now for company B:\[t_b=28x+15\] where \[t_b\] are total cleaning charges for company B moving onto 2nd part of your question for 4 mattresses insert x=4 in both eqns \[t_a=30.4+13=120+13=133\] \[t_b=28*4+15=127\] obviously company B is charging less for 4 mattresses for 3rd part again put x=7 in both eqns \[30*7+13=210+13=223=t_a\] \[28*7+15=211=t_b\] so net saving would be \[223-211=$12\]

OpenStudy (aurorab):

and question, what part of ll this is the seperate parts in the question i origanally asked?

OpenStudy (aurorab):

evermind!! I found em! lol